OpenStudy (anonymous):

Integrate from 0 to 1 (x^3-4x-10)/(x^2-x-6) when I done integrating I get ln|x-3|+2ln|x+2| I keep checking online it says that there is something missing.

OpenStudy (lgbasallote):

are you asking about \[\int \frac{x^3 - 4x - 10}{x^2 - x - 6}dx\]

OpenStudy (anonymous):

yes that's it

OpenStudy (lgbasallote):

well if i were you i'd divide first  so should be \[\large \int (x + 1 + \frac{3x + 4}{x^2 - x - 6})dx\] did you do this?